If the perimiter of a square equals one foot, then what is the area of this square in square inches?
Answer by dkppathak(439)

If the perimiter of a square equals one foot, then what is the area of this square in square inches?
perimeter of square=12 ( we know 1 foot is equal to 12 inches )
let side is a than 4a=12
a=3inches
area =side square
area =3^2=3x3=9 square inch